Evaluate the following polynomial expression using Horner’s rule. Assume the value of x

is 3.

𝑃(𝑥) = 10x5 + 6x4+ 3x3 − 6x2 + 8𝑥 + 15

Show all the intermediate steps.


1
Expert's answer
2021-05-16T21:56:01-0400

"\text{Given, the polynomial }p(x)=10x^5+6x^4+3x^3-6x^2+8x+15 and x_0=3\newline
\text{Since, it is 5 degree polynomial.}
n=5\newline
Horner's rule,\newline
1. Set k = n\newline
2. Let b_k = a_k\newline
3. Let b_{k - 1} = a_{k - 1 }+ b_kx_0\newline
4. Let k = k - 1\newline
5. If k ≥ 0 \text{then go to step 3}\newline
\text{Else End}\newline
Then,\newline
b_5=a_5=10\newline
b_4 = a_4 + x_0b_5=6+3×10=36
\newline
b_3 = a_3 + x_0b_4=3+3×36=111
\newline
b_2 = a_2 + x_0b_3=-6+3×111=327
\newline
b_1 = a_1 + x_0b_2=8+3×327=989
\newline
b_0 = a_0 + x_0b_1=15+3×989=2982
\newline
Therefore,P(3)=2982."